[figure: 9.7 cm]

CORRELATION FUNCTIONS IN THE SPIN-FERMION MODEL

Jan Bała and Andrzej M. Oleś

Institute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland

We investigate a spin-fermion multiband model [1] derived for the valence electrons in CuO_{2} planes including both p_{σ} and p_{π} in-plane oxygen orbitals in the limit of strong correlations. Treating the excitations of the spins at Cu sites within the linear spin-wave theory and solving the problem in a self-consistent Born approximation we have found the sector of Zhang-Rice (ZR) quasiparticles coexisting with higher energy oxygen states. The quasiparticle dispersion and the dispersion of the oxygen states at higher energy are even in quantitative agreement with the experimental data in insulating Sr_{2}CuO_{2}Cl_{2} [2], reproducing the ZR singlet states with their asymmetry in the intensity around the (π/2, π/2) point in the Brillouin zone and giving strong O(2p_{π}) band states near the (π, π) point. Next, using the realistic parameters for CuO_{2} planes we calculated various spin-correlation functions relative to the position of the moving oxygen hole in the quantum Néel state produced by copper spins. The antiferromagnetic correlations are reduced in the neighborhood of the hole and the correlation functions can exhibit similar dipolar distortions as reported in the context of the t-J model [3]. The spin deviations created by the hole motion are weaker than those obtained in the t-J model.

[1] J. Bała, A.M. Oleś, and J. Zaanen, Phys. Rev. B. 54, 10 161 (1996).

[2] J. J. M. Pothuizen et al., Phys. Rev. Lett. 78, 717 (1997).

[3] A. Ramšak and P. Horsch, Phys. Rev. B 48, 10 559 (1993), ibid. 57, 4308 (1998).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Jan Bała

Address for correspondence : Institute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland

Email address : ufbala@jetta.if.uj.edu